OpenStudy (anonymous):

find lcm of 8 and 12

OpenStudy (anonymous):

2

OpenStudy (anonymous):

96

jimthompson5910 (jim_thompson5910):

8: 2*2*2 12: 2*2*3 LCM = 2*2*2*3 = 8*3 = 24 So the LCM is 24

OpenStudy (anonymous):

whoops. I take my medal back

OpenStudy (anonymous):

lol ! i got the wrong answer sry

OpenStudy (anonymous):

it alright =D

jimthompson5910 (jim_thompson5910):

@blue, you can divide that by the GCF 4 to get the answer 24

OpenStudy (anonymous):

yeah, :)

jimthompson5910 (jim_thompson5910):

LCM of x and y = (x*y)/(GCF of x and y)